### v1

# MIBench: Evaluating Multimodal Large Language Models over Multiple Images

Abstract
Built on the power of LLMs, numerous multimodal large language models (MLLMs) have recently achieved remarkable performance on various vision-language tasks across multiple benchmarks. However, most existing MLLMs and benchmarks primarily focus on single-image input scenarios, leaving the performance of MLLMs when handling realistic multiple images underexplored. Although a few benchmarks consider multiple images, their evaluation dimensions and samples are very limited. Therefore, in this paper, we propose a new benchmark MIBench, to comprehensively evaluate fine-grained abilities of MLLMs in multi-image scenarios. Specifically, MIBench categorizes the multi-image abilities into three scenarios: multi-image instruction (MII), multimodal knowledge-seeking (MKS) and multimodal in-context learning (MIC), and constructs 13 tasks with a total of 13K annotated samples. During data construction, for MII and MKS, we extract correct options from manual annotations and create challenging distractors to obtain multiple-choice questions. For MIC, to enable an in-depth evaluation, we set four sub-tasks and transform the original datasets into in-context learning formats. We evaluate several open-source MLLMs and close-source MLLMs on the proposed MIBench. The results reveal that although current models excel in single-image tasks, they exhibit significant shortcomings when faced with multi-image inputs, such as confused fine-grained perception, limited multi-image reasoning, and unstable in-context learning. The annotated data in MIBench is available at https://huggingface.co/datasets/StarBottle/MIBench.
MIBench: Evaluating Multimodal Large Language Models
over Multiple Images
Haowei Liu1,2, Xi Zhang3, Haiyang Xu3†, Yaya Shi4, Chaoya Jiang5, Ming Yan3, Ji Zhang3, Fei Huang3, Chunfeng Yuan1,2†, Bing Li1,2, Weiming Hu1,6 1MAIS, Institute of Automation, Chinese Academy of Sciences, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, China 3Alibaba Group 4University of Science and Technology of China 5Peking University 6School of Information Science and Technology, ShanghaiTech University, China liuhaowei2019@ia.ac.cn, cfyuan@nlpr.ia.ac.cn {shuofeng.xhy, ym119608}@alibaba-inc.com
footnotetext: Corresponding authors.
1 Introduction
Figure 1: Overview of our MIBench, which covers three multi-image scenarios and a total of 13 tasks.
Recently, leveraging the powerful comprehension and reasoning abilities of LLMs, many MLLMs (such as LLaVA-1.5(Liu et al., 2024) and mPLUG-Owl2(Ye et al., 2024)) have demonstrated outstanding performance across various vision-language tasks (e.g. image captioning, VQA and visual grounding). Concurrently, numerous benchmarks (e.g. MME(Fu et al., 2023), MMBench(Liu et al., 2023) and SEED-Bench(Li et al., 2024)) have been proposed to evaluate the abilities of MLLMs in terms of different perspectives such as recognition, localization and reasoning.
However, most existing MLLMs focus on single-image scenarios. Accordingly, previous benchmarks primarily evaluate MLLMs based on single-image inputs. In contrast, real-world multimedia information, such as web pages and social media, generally contains multiple images and corresponding text in interleaved forms. Therefore, multi-image scenarios have greater practical value than single-image scenarios, which makes it worth exploring whether existing single-image MLLMs possess emergent abilities for multi-image inputs. Moreover, some methods like Sparkles (Huang et al., 2023) and Mantis (Jiang et al., 2024) explore multi-image scenarios but have not comprehensively evaluated the models’ multi-image abilities. As shown in Table 1, Sparkles evaluates the model solely on a small-scale multi-image chat dataset, and the assessment relies entirely on scoring by GPT-4. Mantis-Eval focuses on multi-image reasoning and has a limited scale of 217 samples.
In this paper, to comprehensively evaluate the multi-image ability of MLLMs, we propose a large-scale multi-image benchmark MIBench, which covers 13 different tasks with a total of 13K high-quality samples. As shown in Figure 1, MIBench contains three multi-image scenarios, i.e. Multi-Image Instruction (MII), Multimodal Knowledge-Seeking (MKS) and Multimodal In-Context Learning (MIC). MII is a basic multi-image scenario, where the instructions involve perception, comparison and reasoning across multiple images. MKS presents a different scenario, in which models are provided with interleaved image-text data as external knowledge, while the question itself is about a single image or even independent of any image. MIC is another scenario where MLLMs respond to queries (e.g. image & question) by conditioning on a series of multimodal demos. The three scenarios are further divided into 13 different tasks, and examples are shown in Figure 2.
The MII and MKS scenarios comprise 9K multiple-choice questions. To get these questions, we first sample images from nine existing datasets, and convert the original annotations into questions and ground truth options according to the tasks. To obtain challenging distractors and mitigate inherent biases of options, we devise task-specific strategies to sample from annotations or generate distractors using GPT-4. For MKS, we also devise corresponding strategies to sample images and associated texts from the datasets as external knowledge. The MIC scenario contains 4K short-answer questions, covering close-ended VQA, open-ended VQA, object hallucination, and demo-based task learning. We convert the data sampled from four datasets into the VQA format, and retrieve samples of the same task to construct demos. To ensure high quality, we combine automated filtering and manual verification to remove samples with ambiguous or duplicate options. For multiple-choice questions, we use accuracy as the metric and employ circular evaluation (Liu et al., 2023) to mitigate the position bias of LLMs. For short-answer questions, we use exact matching as the metric.
We evaluate several existing MLLMs on the proposed MIBench, including both closed-source (e.g. GPT-4V) and open-source models (e.g. LLaVA-1.5, Qwen-VL and IDEFICS2). The evaluation results reveal that current MLLMs especially open-source models have major flaws in multi-image scenarios. We’ll release our benchmark to spur progress in improving the multi-image abilities of MLLMs.
Our contributions are summarized as follows:
  • We propose the first large-scale and comprehensive benchmark MIBench for evaluating the multi-image abilities of MLLMs, covering three scenarios and 13 tasks.
  • The evaluation on MIBench reveals that existing MLLMs especially open-source models face significant challenges in fine-grained perception and multi-image reasoning.
  • Current MLLMs perform poorly in the multimodal knowledge-seeking scenario, due to limited abilities of multi-image perception and comparison. And there remains considerable room for improvement in multimodal in-context learning abilities.
Benchmark Scenario #Multi-Image Task #Multi-Image Sample Answer Type Evaluator
MME Single-Image 0 0 Yes/No Metrics
MMBench Single-Image 0 0 Multi-choice GPT
SEED-Bench Single-Image 4 829 Multi-choice Metrics
Sparkles-Eval Multi-Image Dialogue 1 150 Open-ended GPT-4
Mantis-Eval Multi-Image Reasoning 1 217 Multi-choice & Short Answer Metrics
MIBench Comprehensive Multi-Image 13 13K Multi-choice & Short Answer Metrics
Table 1: Comparison of the proposed MIBench with recent MLLM benchmarks.
2 Related Work
2.1 Multimodal Large Language Models
Recent research (Zhu et al., 2023; Liu et al., 2024; Dai et al., 2024; Ye et al., 2023) has expanded LLMs (e.g. LLaMA Touvron et al., 2023) into multimodal scenarios, enabling them to process both visual and textual information. Some studies (Jiang et al., 2024; Huang et al., 2023; Laurençon et al., 2024) have further explored augmenting MLLMs with multi-image understanding ability. However, there lacks a comprehensive benchmark for multi-image understanding abilities, which limits the full exploration of these models’ potential and hinders the development of this field.
2.2 MLLM Benchmarks
The rapid development of MLLMs has led to the emergence of a series of benchmarks, like LVLM-eHub (Xu et al., 2023), SEED-bench (Li et al., 2023a), MMBench (Liu et al., 2023), and MM-Vet (Yu et al., 2023). However, these benchmarks primarily focus on single-image evaluation, and often overlook multi-image perception and reasoning abilities, which hold greater practical value.
Recent studies have developed benchmarks for assessing multi-image reasoning capabilities. SparklesEval aims to establish a benchmark for multi-turn dialogues and multi-image scenarios. However, it exclusively focuses on dialogue scenario, and relies entirely on GPT-4 for evaluation. Besides, its scale is small. Other datasets such as Mantis-Eval (Jiang et al., 2024) and SEED-Bench2 (Li et al., 2024) also cover a limited number of multi-image tasks and their scale is constrained due to reliance on manual annotations.
In this paper, we propose a large-scale benchmark with 3 scenarios and 13 tasks, comprising 13K high-quality annotations, to comprehensively evaluate the multi-image capabilities of MLLMs.
3 MIBench
We categorize the multi-image inputs into three scenarios: Multi-Image Instruction (MII), Multimodal Knowledge-Seeking (MKS) and Multimodal In-Context Learning (MIC). As Figure 2 shows, MII refers to cases where instructions involve perception, comparison and reasoning across multiple images. For instance, "Do the two images show the same number of cats?" In contrast, MKS provides models with interleaved image-text data as external knowledge, while the question itself is about an image or independent of any image. MIC is another popular scenario, in which MLLMs respond to visual questions while being provided with a series of multimodal demonstrations (i.e., demos).
3.1 Evaluation Taxonomy
Figure 2: Examples of the multi-image scenarios with a total of 13 tasks. The correct answers are marked in blue.
3.1.1 Multi-Image Instruction
General Comparison (GC) task examines the model’s general understanding of each image (e.g. scene, attribute and location), and comparison across different images. GC represents the most fundamental aspect of multi-image abilities. We use the image-pair description dataset NLVR2 (Suhr et al., 2018) for data construction.
Subtle Difference (SD) task examines the model’s ability to perceive subtle differences between similar images. Compared to general comparison, the SD task requires more fine-grained perception ability. The image editing dataset MagicBrush (Zhang et al., 2024) is adopted in this task.
Visual Referring (VR) task evaluates whether the model can utilize the referring information provided by input images to comprehend the relationships between different objects. Figure 2(c) shows an example of the VR task, whose data is from the visual relation dataset VrR-VG (Liang et al., 2019).
Temporal Reasoning (TR) task assesses the model’s understanding of the temporal relationships among a series of consecutive images, and its comprehension of the overall content conveyed by these images. We employ the video understanding dataset Something-Something V2 (Goyal et al., 2017a) for this task.
Logical Reasoning (LR) task requires the model to perform logical reasoning and analyze the causal relationships between objects or events shown in the images. The video QA dataset NExT-QA (Xiao et al., 2021) is used for data construction.
3.1.2 Multimodal Knowledge-Seeking
Fine-Grained Visual Recognition (FVR) task examines the model’s ability to recognize the object in the query image when given multiple reference images. It requires the model to understand the image-label correspondence in the reference images, as well as link similar images. A combination of several fine-grained recognition datasets (Khosla et al., 2011, Wah et al., 2011 and Nilsback and Zisserman, 2008) is used for this task.
Text-Rich Images (TRI) VQA task evaluates the model’s ability to understand text-rich images and extract information relevant to the given question, which is very common in real-world scenarios (e.g. reading slides or documents). We adopt the SlideVQA (Tanaka et al., 2023) dataset for data construction.
Vision-linked Textual Knowledge (VTK) task corresponds to a very practical scenario where the question is beyond the visual content of the query image, such as querying background knowledge. The provided outside knowledge encompasses images and corresponding text possibly retrieved from a knowledge base (e.g., Wikipedia). The model is required to link the query image to the relevant image, and extract useful information from the corresponding text. Figure 2(h) shows an example, whose data is from the InfoSeek dataset (Chen et al., 2023).
Text-linked Visual Knowledge (TVK) task refers to cases where the text-only question is about the visual attributes of a specific object. For instance, "Is the China National Stadium round or square?". When provided with outside knowledge in an interleaved image-text form, the model needs to link the question to the relevant text, and extract visual information from the corresponding image. This task is very common in real life such as browsing web pages. Figure 2(i) shows an example, whose data is from the WebQA dataset (Chang et al., 2022).
3.1.3 Multimodal In-Context Learning
The in-context learning ability enables LLMs to gain performance boost when provided with a series of demos. Flamingo (Alayrac et al., 2022), OpenFlamingo (Awadalla et al., 2023) and IDEFICS (Laurençon et al., 2024) have explored multimodal in-context learning (MIC). For the evaluation of the MIC ability, existing methods solely assess the model’s performance via a holistic metric, such as accuracy on the VQAv2 (Goyal et al., 2017b) dataset. To evaluate the model’s MIC ability in a fine-grained manner, we categorize the MIC scenario into the following four distinct tasks:
Close-ended VQA task requires the model to answer from a predefined set of responses, which is provided via multimodal demos. This task examines the model’s ability to learn the image-label mapping relationships from the demos. We use the Mini-ImageNet dataset (Vinyals et al., 2016) for data construction.
Open-ended VQA task has an open range of possible answers which cannot be fully covered by the demos. The task evaluates the model’s ability to learn task patterns through demos. We conduct a balanced sampling of different knowledge types of the OK-VQA dataset (Marino et al., 2019) for this task.
Hallucination phenomenon is a significant challenge faced by MLLMs. In this task, we construct the hallucination dataset POPE (Li et al., 2023b) into in-context learning format, and study the impact of the model’s MIC ability on the hallucination phenomenon.
Demo-based Task Learning is a core aspect of in-context learning, which enables the model to rapidly adapt to new tasks given a few demos. To investigate existing MLLMs’ demo-based task learning ability, we select several visual tasks from the VQAv2 dataset and remove the task instructions. Instead, we present the task demos in the form like "rabbit: 3". Figure 2(m) shows an example.
3.2 Data Generation
In Section 3.1, we introduced the evaluation tasks and the corresponding data source of the proposed MIBench. However, the generation of test samples using the original annotations is nontrivial. We meticulously devise a data generation pipeline, including various strategies of question generation, distractor generation and external knowledge sampling for different tasks.
Question Generation. To enhance the diversity of questions, we devise corresponding prompts for each task, and employ GPT-4 to generate a variety of question forms. We then randomly sample from the question pool to construct the test samples. For instance, for the General Comparison (GC) task, the questions such as "Is the subsequent sentence an accurate portrayal of the two images?" and "Can the given sentence accurately illustrate what’s in these two images?" are utilized.
Distractor Generation. For different tasks, we adopt two methods of distractor generation. One way is to sample from the original annotations following certain strategies. For instance, for the Temporal Reasoning (TR) task, we utilize the Something-something V2 dataset for data construction. To prevent the model from taking shortcuts by identifying objects to choose the correct options, we sample different temporal relationships of the same object from the annotations as distractors. In this way, the constructed test samples can more accurately reflect the model’s understanding of temporal relationships. The second method is to generate distractors with the help of GPT-4. For instance, in the Text-Rich Images (TRI) task, we prompt GPT-4 to generate distractors according to the question and the correct answer.
External Knowledge Sampling. For the Multimodal Knowledge-Seeking (MKS) scenario, reasonably sampling interleaved image-text data as external knowledge is very important to the quality of test samples. For instance, in the Vision-linked Textual Knowledge (VTK) task, we select text and images from the original annotations which have the same question as the current query but with different answers as external knowledge. This approach avoids selecting text and images unrelated to the current query, and can thus generate more challenging distractors.
3.3 Quality Control
To mitigate data contamination, our construction of test data exclusively utilizes the validation or test sets from existing datasets. Furthermore, to ensure the high quality and reliability of the test data, we combine automated filtering and manual verification to remove samples with ambiguous or duplicate options.
Moreover, in the Multimodal Knowledge-Seeking scenario, our goal is to assess the model’s ability to utilize external knowledge to answer questions, rather than relying on its inherent knowledge. Therefore, to prevent the varying inherent knowledge of different models from impairing the evaluation, we conduct pre-tests using multiple MLLMs and eliminate those samples which the models can correctly answer without external knowledge.
3.4 Evaluation
For the multiple-choice questions, we employ the accuracy of the predicted options as the evaluation metric. Notably, early MLLMs such as mPLUG-Owl tend to produce longer responses rather than directly outputting the option. For these models, we use GPT-4 to determine which option matches the predicted content. In addition, similar to the observation of MMBench, we find that different MLLMs show preferences for specific options (i.e. position bias). Therefore, we set the correct option sequentially to ’A’, ’B’, ’C’ and ’D’. A model is considered to have correctly answered a sample only if it consistently provides the correct response across multiple tests. In this way, the impact of position bias on the evaluation results is mitigated.
4 Experiments
Model Multi-Image Instruction Multimodal Knowledge-Seeking
GC SD VR TR LR FVR TRI VTK TVK
Closed-source MLLMs
GPT-4o 80.7 90.5 46.8 68.0 69.8 98.3 74.8 54.7 63.3
GPT-4V 72.8 79.2 45.8 61.8 66.3 90.2 71.0 52.0 56.0
Open-source MLLMs
Mantis 83.0 54.1 37.6 45.5 63.4 16.4 37.7 26.4 41.7
IDEFICS2-I 83.1 49.7 32.6 44.8 56.4 42.4 43.9 25.6 39.0
MMICL 53.7 46.4 41.1 47.0 59.6 56.6 27.6 22.1 35.9
mPLUG-Owl2 64.2 40.1 35.6 30.7 41.3 13.3 39.0 17.0 25.6
Qwen-VL 45.9 22.5 16.3 27.5 36.8 58.8 35.9 22.9 18.1
LLaVA-1.5 40.6 14.9 24.1 30.1 44.8 18.2 25.8 16.7 26.3
mPLUG-Owl 19.1 4.0 21.7 8.0 29.2 17.3 12.1 14.9 20.6
Table 2: Evaluation results on the multi-image instruction and multimodal knowledge-seeking scenarios of MIBench.
4.1 Models
In this section, we evaluate MLLMs using the constructed MIBench dataset. We first evaluate MLLMs on the Multi-Image INstruction and Multimodal Knowledge-Seeking scenarios. These models can be categorized into three distinct groups: (1) close-source models, including GPT-4V and GPT-4o; (2) open-source single-image MLLMs, including mPLUG-Owl, LLaVA-1.5, Qwen-VL and mPLUG-Owl2; (3) open-source models natively supporting multi-image input, including Mantis and IDEFICS2. For the open-source models, we employ greedy decoding for prediction generation.
Then we evaluate MLLMs on the Multimodal In-Context Learning scenario. Currently, apart from OpenFlamingo, MMICL and IDEFICS, most MLLMs have not explored multimodal ICL abilities. As a result, they exhibit consistent negative ICL effect in testing, i.e. the increase of demos leads to performance decline. Therefore, we only present the results of models that possess multimodal ICL abilities.
4.2 Evaluation Results
Figure 3: A qualitative case of Subtle Difference task.
Figure 4: Evaluation results on the Multimodal In-Context Learning scenario.
4.2.1 Multi-Image Instruction & Multimodal Knowledge-Seeking
Table 2 shows the main results of the Multi-Image Instruction (MII) and Multimodal Knowledge-Seeking (MKS) scenarios. Through these results, we have several valuable observations:
Close-source MLLMs exhibit superior performance than close-source models. In most MII and MKS tasks, the performance of open-source models lags significantly behind that of proprietary models. For instance, on the Subtle Difference (SD) task, the best-performing open-source model Mantis achieves an accuracy of 54.1%, falling behind GPT-4o by 36.4%.
Open-source MLLMs are inadequate in fine-grained perception tasks. Although many open-source MLLMs have decent performance on the General Comparison (GC) task, their performance is significantly worse on the Subtle Difference (SD) task. For instance, IDEFICS2-I achieves 83.1% and 49.7% on the GC and SD tasks respectively. In contrast, GPT-4V and GPT-4o largely outperform open-source models, due to their high-resolution input strategy. Furthermore, Figure 3 provides a qualitative case supporting the above point.
Multi-image MLLMs perform better than single-image models on most tasks. This verifies that pre-training on interleaved image-text data (e.g. IDEFICS2) and instruction tuning on multi-image data (e.g. Mantis) are beneficial for improving multi-image abilities.
The Visual Referring (VR) task is particularly challenging for existing MLLMs. As it requires integration of fine-grained perception, spatial correspondence and relation reasoning, most models have not achieved satisfactory performance on the VR task. Even the best-performing model GPT-4o has not exceeded a 50% accuracy rate.
Most existing open-source MLLMs perform poorly in the Multimodal Knowledge-Seeking (MKS) scenario. For instance, the performance of mPLUG-Owl2 in both categories of Vision-linked Textual Knowledge (VTK) and Text-linked Visual Knowledge (TVK) is below 30%. In the Fine-grained Visual Recognition (FVR) task, which requires the combination of fine-grained perception and comparison abilities, mPLUG-Owl2’s performance is merely over 10%. Compared to single-image MLLMs, multi-image models such as IDEFICS2-I show better capabilities in utilizing multimodal external knowledge. However, there is still significant room for improvement, as the performance of IDEFICS2-I on both VTK and TVK tasks is under 40%.
4.2.2 Multimodal In-Context Learning
[…]
The current models exhibit multimodal ICL abilities on close-ended VQA. As shown in Figure 4(a), after removing the images in the demos, the performance of most models declines, and the extent of this decline increases with the number of shots. This indicates that these models have learned the image-label mapping relationships in the demos, exhibiting a certain degree of multimodal ICL ability.
Multimodal ICL abilities of different models appears to be driven by different modalities. As shown in Figure 4(b), when the number of demos increases, all models show consistent performance improvement. However, for OpenFlamingo, removing images from the demos does not cause a significant performance change, indicating that OpenFlamingo’s ICL on this task is primarily driven by text. In contrast, the absence of images leads to a significant performance decline for IDEFICS and MMICL, indicating that they possess a certain degree of multimodal ICL ability.
Multimodal ICL abilities of current models do not alleviate the hallucination phenomenon. As shown in Figure 4(c), on object hallucination task, only IDEFICS-I and IDEFICS2-I exhibit slight accuracy improvements with an increasing number of shots, while other models show negative effects. It indicates that ICL provides very limited help in mitigating hallucinations and may even exacerbate them. Comparing the base and instruction-tuned versions of IDEFICS, it is evident that instruction tuning can help alleviate hallucinations.
Most models possess some capacity of demo-based task learning, but the capacity is relatively limited. Figure 4(d) shows the model performance under different shots in counting and color tasks demonstrated only through examples. It is evident that with an increasing number of demos, IDEFICS shows significant gains, OpenFlamingo quickly reaches saturation, and MMICL even fails to follow the task format presented in the demos. In fact, except for MMICL, these models can effectively follow the output format with just 4 shots, and their performance improves with more shots. It reflects that OpenFlamingo and IDEFICS possess a certain degree of demo-based task learning ability. In addition, compared to the experimental results with explicit task instructions (e.g., ‘How many people are in the room?’), there remains a significant performance gap, indicating that the demo-based task learning abilities of current models still have substantial room for improvement.
4.3 Analysis
Text-rich Images VQA Text-linked Visual Knowledge Vision-linked Textual Knowledge
w/ Dis. w/o Dis. w/ Dis. w/o Dis. w/ Dis. w/o Dis.
mPLUG-Owl2 39.0 42.1 25.6 29.6 17.0 90.1
Mantis 37.7 42.6 41.7 47.7 26.4 88.1
IDEFICS2-${\mathrm{I}}$ 43.9 46.8 (59.5) 39.0 45.2 25.6 91.0
Table 3: Ablation study of the impact of distractors on various tasks on the multimodal knowledge-seeking scenario.
Confusion Reasoning
Conf. A Conf. B Tem. Obj.
mPLUG–Owl2 87.0 25.0 30.7 56.6
Qwen–VL 89.2 26.8 27.5 60.9
LLaVA–1.5 91.8 31.6 30.1 59.3
Mantis 91.2 83.6 45.5 75.7
Table 4: Ablation study on the multi-image confusion phenomenon and the temporal reasoning task.
4.3.1 Multi-image Confusion Phenomenon
When evaluating MLLMs on the MIBench benchmark, we observe that open-source MLLMs, particularly single-image models, exhibit confusion when handling multiple images. To validate this issue, we derive two confusion subsets with 500 samples respectively from the POPE dataset used in the hallucination task. In subset A, each sample consists of one image and one question. The question asks whether a specific object is present in the image, which actually is not contained. In subset B, an extra image containing the object in the question is added to each sample in subset A as a distractor. As shown in Table 4, it can be observed that the performance of the three single-image models significantly decline after the addition of the extra image, while the multi-image model Mantis also has a slight performance drop. It confirms that current open-source MLLMs, especially single-image models, suffer from severe confusion, thereby affecting their performance in multi-image instructions and multimodal knowledge-seeking.
4.3.2 Limited Reasoning Ability
In the construction of temporal reasoning, we utilize the ground truth of videos as the correct option and sample different actions of the same object as distractors. Under this setting, the majority of MLLMs demonstrate poor performance. To further study these results, we replace the same objects in the distractors with different objects and test several representative models. As indicated in the table, under the setting where distractors contain different objects, the performance of mPLUG-Owl2, LLaVA-1.5 and Mantis models significantly improves, as the models can take shortcuts by distinguishing between objects. The results indicate that for current MLLMs, the reasoning ability across multiple images is significantly inferior to their spatial perception ability.
4.3.3 Bottlenecks of the MKS task
Compared to multi-image instruction, multimodal knowledge-seeking requires the model to extract relevant information from external image-text knowledge sources and is thus more challenging. To investigate the bottlenecks of multimodal knowledge-seeking tasks, we compare the impact of distracting content.
As shown in Table 3, for text-linked visual knowledge, removing distracting content(i.e., only retaining the information relevant to the question) results in a certain performance improvement. It indicates that the model’s ability to identify relevant information (i.e., link by text) still can be improved. On the other hand, even after the removal of distracting content, the performance remains poor. It suggests that the primary bottleneck for this task is the deficiencies of MLLMs in perceiving and reasoning with visual information.
In contrast, for the task of vision-linked textual knowledge, the removal of distracting content leads to a significant performance improvement. It suggests that the bottleneck for this task lies in the MLLMs’ ability to mine effective messages through image comparison (i.e., link by image).
On text-rich images VQA, removing distracting content brings some performance boost. Based on this, IDEFICS2-I further boosts from 46.8% to 59.5% by employing image splitting for higher resolution inputs. The significant performance gain indicates that the bottleneck of this task is more related to information loss caused by low resolution.
From the above comparisons, it can be concluded that the current MLLMs’ abilities in perceiving, contrasting, and reasoning with visual information are remarkably inferior to their abilities with text, and face substantial challenges in understanding rich-text images due to resolution issues.
5 Conclusion
While MLLMs have shown strong performance in various vision-language tasks, their abilities with multi-image inputs remain underexplored. To address this gap, we introduce MIBench in this paper, a benchmark that evaluates MLLMs in three multi-image scenarios: multi-image instruction, multimodal knowledge-seeking, and multimodal in-context learning, across 13 tasks with 13K annotated samples. The evaluation results reveal that existing models, despite excelling in single-image tasks, face significant challenges with multi-image inputs. We will release our benchmark’s annotated data to facilitate further research. We hope this work spurs progress in improving the multi-image abilities of MLLMs.
6 Limitations
Due to the input length limitations of current MLLMs, our constructed benchmark includes 2 to 8 input images in each sample. However, real-world scenarios may involve a large number of images. We’ll investigate the evaluation with more images in future work.